# Toricity in families of Fano varieties

### v1

@@ 6. Examples and questions
In this section, we collect some examples of the behavior of toricity and cluster type in families of Fano varieties. We propose some questions for further research. The first two exmaples show that the property of being a toric variety is neither open nor closed.
@@
The projective plane $\mathbb{P}^{2}$ admits non-toric, quasismooth degenerations. Namely, let $(a,b,c)$ and $(a,b,d)$ be two pairs of Markov triples for which $d=3ab-c$, the so-called adjacent Markov triples. Consider the weighted hypersurface
@@
By [28], this is a non-toric, quasismooth, rational, Fano $\mathbb{G}_{m}$-surface of Picard rank one, which is a degeneration of ${\mathbb{P}}^{2}$. Furthermore, by [28, Theorems 1.2 and 1.3], all the non-toric, normal, rational projective $\mathbb{G}_{m}$-surfaces which are degenerations of ${\mathbb{P}}^{2}$ arise in this way. Hence, the property of being a toric variety is not closed in families of Fano varieties.
@@
it has $10$ nodal singularities and its automorphism group is $S_{6}$; in particular, $X$ is not a toric variety. It remains to construct a boundary $B$ on $X$ such that $(X,B)$ is a log CY pair of cluster type. There is a small resolution $\pi\colon Y\rightarrow X$ which is isomorphic to the blow-up $\phi\colon Y\to{\mathbb{P}}^{3}$ at $5$ general points $\{p,q,r,s,t\}\subset\mathbb{P}^{3}$ [29]. Let $H_{1}$ (resp. $H_{2}$) by the hyperplane in ${\mathbb{P}}^{3}$ spanned by $\{p,q,r\}$ (resp. $\{p,s,t\}$). Since $\{p,q,r,s,t\}$ are in general position, the line $\ell_{1}\coloneqq H_{1}\cap H_{2}$ only contains one point of the set $\{p,q,r,s,t\}$, namely $p$. Next, let $u_{1}$ and $u_{2}$ be two general points in $\ell_{1}$, and let $H_{3}$ (resp. $H_{4}$) be the hyperplane in ${\mathbb{P}}^{3}$ spanned by $\{u_{1},q,s\}$ (resp. $\{u_{2},r,t\}$). By construction, the pair $({\mathbb{P}}^{3},H_{1}+H_{2}+H_{3}+H_{4})$ is log Calabai–Yau, and each point of the set $\{p,q,r,s,t\}$ is contained in precisely two of the hyperplanes $\{H_{1},H_{2},H_{3},H_{4}\}$. In particular, if we write
@@
Let $X_{0}$ be the Gorenstein del Pezzo surface of Picard rank one with a single $D_{5}$ singularity. By [23, Theorem 1.1 and Example 8.3], we know that $X_{0}$ admits a ${\mathbb{Q}}$-Gorenstein smoothing $\mathcal{X}\rightarrow\mathbb{A}^{1}$. Since the anticanonical volume of $X_{0}$ is $4$ (see, e.g., [39, Lemma 3]), the general fiber of $\mathcal{X}\rightarrow\mathbb{A}^{1}$ is a smooth del Pezzo surface of degree 5. Every smooth del Pezzo surface of degree 5 admits a $1$-complement of coregularity zero [1, Theorem 2.1] and so it is of cluster type by [22, Lemma 1.13]. On the other hand, the central fiber $\mathcal{X}_{0}$ is not of cluster type, as cluster type surfaces only admit $A$-type singularities (see [15, Proposition 5.1]).
@@
Consider the trivial family $\pi_{2}\colon\mathcal{Y}\coloneqq{\mathbb{P}}^{2}\times\mathbb{A}^{1}_{t}\rightarrow\mathbb{A}^{1}_{t}$. Now, let $\mathcal{D}$ be a divisor on $\mathcal{Y}$ for which $\mathcal{D}_{0}$ is the reduced sum of the coordinate hyperplanes and $\mathcal{D}_{t}$ is a smooth cubic curve for every $t\neq 0$. Let $\mathcal{X}$ be the deformation of $\mathcal{Y}$ to the normal cone of $\mathcal{D}$. For each $t\in\mathbb{A}^{1}$, we let $\psi_{t}\colon\mathcal{X}_{t}\rightarrow\mathbb{A}^{1}_{s}$ be the induced deformation of $\mathcal{Y}_{t}$ to the normal cone over $\mathcal{D}_{t}$. Then, the threefold $\mathcal{X}$ admits a morphism $\psi\colon\mathcal{X}\rightarrow\mathbb{A}^{2}$ whose fiber over $(s,t)$ is $\psi_{t}^{-1}(s)$. By construction, all the fibers of $\psi$ over $\mathbb{G}_{m}\times\mathbb{A}^{1}_{t}$ are isomorphic to ${\mathbb{P}}^{2}$ and hence toric varieties. The fibers over $\{0\}\times\mathbb{G}_{m}$ are cones over smooth cubic curves and so they are not toric. The fiber over $(0,0)$ is the non-normal cone over the reduced sum of the hyperplane coordinates of ${\mathbb{P}}^{2}$, which is a toric variety. We conclude that the subset of $\mathbb{A}^{2}$ parametrizing toric fibers of $\psi\colon\mathcal{X}\rightarrow\mathbb{A}^{2}$ is

### v2

@@ 6. Examples and questions
In this section, we collect some examples of the behavior of toricity and cluster type in families of Fano varieties. We propose some questions for further research. The first two examples show that the property of being a toric variety is neither open nor closed.
@@
The projective plane $\mathbb{P}^{2}$ admits non-toric, quasismooth degenerations. Namely, let $(a,b,c)$ and $(a,b,d)$ be two pairs of Markov triples for which d=3abcd=3ab-c, the so-called adjacent Markov triples. Consider the weighted hypersurface
@@
By [29], this is a non-toric, quasismooth, rational Fano $\mathbb{G}_{m}$-surface of Picard rank one, which is a degeneration of ${\mathbb{P}}^{2}$. Furthermore, by [29, Theorems 1.2 and 1.3], all the non-toric, normal, rational projective $\mathbb{G}_{m}$-surfaces which are degenerations of ${\mathbb{P}}^{2}$ arise in this way. Hence, the property of being a toric variety is not closed in families of Fano varieties.
@@
it has $10$ nodal singularities and its automorphism group is $S_{6}$; in particular, $X$ is not a toric variety. It remains to construct a boundary $B$ on $X$ such that $(X,B)$ is a log CY pair of cluster type. There is a small resolution $\pi\colon Y\rightarrow X$ which is isomorphic to the blow-up $\phi\colon Y\to{\mathbb{P}}^{3}$ at $5$ general points $\{p,q,r,s,t\}\subset\mathbb{P}^{3}$ [30]. Let $H_{1}$ (resp. $H_{2}$) by the hyperplane in ${\mathbb{P}}^{3}$ spanned by $\{p,q,r\}$ (resp. $\{p,s,t\}$). Since $\{p,q,r,s,t\}$ are in general position, the line $\ell_{1}\coloneqq H_{1}\cap H_{2}$ only contains one point of the set $\{p,q,r,s,t\}$, namely $p$. Next, let $u_{1}$ and $u_{2}$ be two general points in $\ell_{1}$, and let $H_{3}$ (resp. $H_{4}$) be the hyperplane in ${\mathbb{P}}^{3}$ spanned by $\{u_{1},q,s\}$ (resp. $\{u_{2},r,t\}$). By construction, the pair $({\mathbb{P}}^{3},H_{1}+H_{2}+H_{3}+H_{4})$ is log Calabi–Yau, and each point of the set $\{p,q,r,s,t\}$ is contained in precisely two of the hyperplanes $\{H_{1},H_{2},H_{3},H_{4}\}$. In particular, if we write
@@
Let $X_{0}$ be the Gorenstein del Pezzo surface of Picard rank one with a single $D_{5}$ singularity. By [24, Theorem 1.1 and Example 8.3], we know that $X_{0}$ admits a $\mathbb{Q}$}-Gorenstein smoothing $\mathcal{X}\rightarrow\mathbb{A}^{1}$. Since the anticanonical volume of $X_{0}$ is $4$ (see, e.g., [41, Lemma 3]), the general fiber of $\mathcal{X}\rightarrow\mathbb{A}^{1}$ is a smooth del Pezzo surface of degree $4$. Every smooth del Pezzo surface of degree $4$ admits a $1$-complement of coregularity zero [1, Theorem 2.1] and so it is of cluster type by [22, Lemma 1.13]. On the other hand, the central fiber $\mathcal{X}_{0}$ is not of cluster type, as cluster type surfaces only admit $A$-type singularities (see [15, Proposition 5.1]).
@@
Consider the trivial family $\pi_{2}\colon\mathcal{Y}\coloneqq{\mathbb{P}}^{2}\times\mathbb{A}^{1}_{t}\rightarrow\mathbb{A}^{1}_{t}$. Now, let $\mathcal{D}$ be a divisor on $\mathcal{Y}$ for which $\mathcal{D}_{0}$ is the reduced sum of the coordinate hyperplanes and $\mathcal{D}_{t}$ is a smooth cubic curve for every $t\neq 0$. Let $\mathcal{X}$ be the deformation of $\mathcal{Y}$ to the normal cone of $\mathcal{D}$. For each $t\in\mathbb{A}^{1}$, we let $\psi_{t}\colon\mathcal{X}_{t}\rightarrow\mathbb{A}^{1}_{s}$ be the induced deformation of $\mathcal{Y}_{t}$ to the normal cone over $\mathcal{D}_{t}$. Then, the threefold $\mathcal{X}$ admits a morphism $\psi\colon\mathcal{X}\rightarrow\mathbb{A}^{2}$ whose fiber over $(s,t)$ is $\psi_{t}^{-1}(s)$. By construction, all the fibers of $\psi$ over $\mathbb{G}_{m}\times\mathbb{A}^{1}_{t}$ are isomorphic to ${\mathbb{P}}^{2}$ and hence toric varieties. The fibers over $\{0\}\times\mathbb{G}_{m}$ are cones over smooth cubic curves and so they are not toric. The fiber over $(0,0)$ is the non-normal cone over the reduced sum of the coordinate hyperplanes of ${\mathbb{P}}^{2}$, which is a toric variety. We conclude that the subset of $\mathbb{A}^{2}$ parametrizing toric fibers of $\psi\colon\mathcal{X}\rightarrow\mathbb{A}^{2}$ is